Project Bronco, Inc., imposes a payback cutoff of three years for its international investment projects.

  

Year Cash Flow (A) Cash Flow (B)
0 –$ 52,000 –$ 62,000
1 19,000 11,000
2 20,000 14,000
3 17,000 18,000
4 4,000 222,000


What is the payback period for both projects? (Round your answers to 2 decimal places, e.g., 32.16.)

Project A ______ years

Project B ______years

What project should the company accept?

a.) project A

b.) project B
Answer #1

a).

Year Cash Flow(A) Cumulative Cash Flow(A) Cash Flow(B) Cumulative Cash Flow(B)
0 -$52,000 -$52,000 -$62,000 -$62,000
1 $19,000 -$33,000 $11,000 -$51,000
2 $20,000 -$13,000 $14,000 -$37,000
3 $17,000 $4,000 $18,000 -$19,000
4 $4,000 $8,000 $222,000 $203,000

Payback Period = Years before full recovery +

[Unrecovered cost at start of the year / CF during the year]

Payback Period(A) = 2 + [$13,000/$17,000] = 2 + 0.76 = 2.76 years

Payback Period(B) = 3 + [$19,000/$222,000] = 3 + 0.09 = 3.09 years

b). Project A should be selected as its payback period is lesser, and is the only project within the specified limit.

As the payback period is the length of time required to recover the cost of an investment.